OpenStudy (brittany):

solve by factoring or quadratic formula x^2-8x+16x=0

OpenStudy (dumbcow):

This is a perfect square. The middle number is double the square root of last number. sqrt(16) = 4 x^2 -8x+16 = (x-4)^2

OpenStudy (brittany):

and then what?

OpenStudy (dumbcow):

wait is that supposed to be "16x" , didn;t see that

OpenStudy (brittany):

no just 16 sorry

OpenStudy (dumbcow):

well now its in factored form (x-4)^2 = 0 solve for x square root both sides

OpenStudy (brittany):

im confused how do u solve tht equation??

OpenStudy (dumbcow):

im telling you how Take square root of both sides what do you get?

OpenStudy (brittany):

i dont know?

OpenStudy (brittany):

shouldnt the factored form be like (x+4)(x+4)

OpenStudy (dumbcow):

yes that is another way of writing it, except it should be (x-4)(x-4)

OpenStudy (brittany):

so then whats the next step after (x-4)(x-4)

OpenStudy (dumbcow):

find a x value that makes (x-4) = 0 x-4 = 0 solve for x

OpenStudy (brittany):

x=4

OpenStudy (dumbcow):

btw \[\sqrt{(x-4)^{2}} = x-4\] that is important to know the square root always cancels out the square

OpenStudy (dumbcow):

yes answer is x=4

OpenStudy (brittany):

thanks

OpenStudy (dumbcow):

your welcome